OpenStudy (anonymous):

The line passes through the points ( 2, 1) and ( 5, 4) a. y = x + 1 c. y = x - 1 b. y = -x - 1 d. y = x + 3

OpenStudy (anonymous):

@Blank 

OpenStudy (anonymous):

y=mx+b form is what you want. So first thing is to find the slope. \[\huge m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

so2

OpenStudy (anonymous):

You can use point - slope form here once you find the slope \[\huge y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

1*

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

the slope is one

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

So now use point - slope form :)

OpenStudy (anonymous):

its c